OpenStudy (anonymous):

what the correct simplification of (x^3/y^5)^2

OpenStudy (anonymous):

Hint \[\LARGE (x^a)^b=x^{a\cdot b}\] in this case you have.. \[\LARGE \left(\frac{x^3}{y^5}\right)^2=\frac{(x^3)^2}{(y^5)^2}=?\]

OpenStudy (anonymous):

@andrea_loxx do you understand it ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Glad to hear that ;)